Convert 1-i / 2 into trigonometric form


1
Expert's answer
2022-06-07T15:30:22-0400
"r^2=x^2+y^2=(1)^2+(-\dfrac{1}{2})^2=\dfrac{5}{4}"

"\tan\theta=\dfrac{y}{x}=\dfrac{-1/2}{1}=-\dfrac{1}{2}"

"z=1-\dfrac{i}{2}"

Quadrant IV


"r=\dfrac{\sqrt{5}}{2}, \theta=2\pi-\tan^{-1}(\dfrac{1}{2})"

or


"r=-\dfrac{\sqrt{5}}{2}, \theta=\pi-\tan^{-1}(\dfrac{1}{2})"